Nominal height of Cerenkov box is 333cm from top of box to bottom of
bracket into which adjustable foot is attached. The foot adds an
This was measured with tape measure by DMW DJL, Jan 2002.
Extra space is required on both sides of box
versus nominal dimensions given in Ed's memory
gas windows to
Jim Kilmer High-rise 11 East
Support Mirrors in Dave Christian's office
(Hirise 10 West towards North)
Steve Pordes (Por-dez) head of PPD could help finding
cables for bases.
Drawings at Nevis: contact Brian Cole, Columbia
| | | |
|W | | E|
| | | |
Summary of May 2002 trip (PMT work)
November 21, 2003 Phototube installation green = type 2 yellow = type 3 Remaing to install (waiting for wavelength coating) 35=3 38=3 37=3 4=2 8=2 29=2 32=2 3=2 7=1 33=2 36=2 2=2 6=1 1=2 5=1 Base coating person: Eileen Hahn 4164
December 15-18, 2003 tube location=base # 35=91 38=85 39=69 28=86 32=94 4=93 8=95 34=10 37=88 3=96 7=45 2=92 6=80 33=89 36=89 1=90 5=63 Spider box ribbon cables SB1 X6-7 SB2 X6-8 SB3 X6-6 SB Spare Top X6-5 SB4 X6-2 SB5 X6-3 SB6 X6-4 SB Spare Bottom X6-1 Minibox low voltage cables MB1 L3 MB2 L4 MB3 L1 MB4 L2 Pierrick, Needs 8 slots in upstream crate Using LeCroy 3377 TDCs (32 channel 0.5 ns resolution Newly Coated Tubes, were never tested 28131 93 (base was formerly labelled #55, but was apparently a duplicate) 27068 85 27067 88 26509 87 28216 not used One can print using: flpr -q e907_hp4si - single sided and flpr -q e907_hp4si_d - double sided printing
January 14-16, 2004 Called Carl Rosenfeld about delay board 250nsec delay completely passive (no power), chip is just a delay line chips are symmetric, so can plug in up-side down Our boards: Lemo input on chip side Ribbon output on chip side lemo output on solder side (just shorted to ribbon output, so lemo will work, but no impedence match) Our boards are different from TOF TOF boards use lemo input (chip side) but split output 90/10 onto two lemos, ribbon ignored Solder Side 2 1 18 17 34 33 4 3 20 19 36-35 ... 16-15 32-31 48-47 Chip Side 33-34 17-18 1-2 35-36 19-20 3-4 ... 47-48 31-32 15-16 Minibox threshold test term (ohm) measured connected 1 16.8 22.7 5.71 MegaOhm (I believe this is correct) 2 25.2 31.2 5.71 MOhm (I believe this is correct) 3 infinity infinity 206 Ohm (problem) 4 52.2 55.9 1066 Ohm (threshold and low voltage location 76 not connected) (problem) use kinit -a dmw to get kerberos tickets that include afs on fnalu use 'setup netscape' to get access to netscape CKOV DAQ (all Lecroy except where noted) ADC 4300 (some 4300B, does it matter)? slots 2,3,4 Disc 4415A, TDC 3377, slots 6,7 9,10 12,13 DAC Bira 5408 slot 15 Output register DSP PR612 slot 16 LV power supply test (supplies were incorrectly labelled! -> fixed by me) Top MB1 -5.63 V (measured at connector on box, no load) Top MB2 -5.65 Bot MB3 -5.71 Bot MB4 -5.73 Gas System 5000 sccm flow controller trip valve settings .25in top 2in bot bubbler 0.5 oil top Minibox test (check resistance across fan-out connector MB3 most bases are 2.9MOhm, some are 1k-11k (I believe this is bad) Low voltage connector, female end A,B=Gnd; C,D=-5.2 V expected (at detector side, under full load)
February 4-6, 2004 Page system works now, press page and then zero Base LV&threshold resistance test Connectors are 25 pair, last pair is not used MB3: Bot (Beam left) pair base location thresh lv # # # ohm ohm 1 41 62 11k 9k 20 8 55 10k 8k Total with pair 1 and 20 disconnected thresh=lv=1.97 Mohm MB4: Bot (Beam right) pair base location thresh lv # # # ohm ohm 2 72 68 8k 7k 4 46 65 53k 56k 7 43 95 10k 15k 9 66 77 2k 200 16 71 72 1.1M 1.1M 18 62 73 5k 3k Total with pairs 2,4,7,9,16,18 discon thresh=lv=2,04 Mohm MB1 Top Beam Left Total=5.60 MOhm pair base loc thresh lv lvconnector pair # # # ohm ohm # # 7 77 40 53k 52k 7bar 19 8* 69 39 open* * = fixed by pushing on harder/twisting total (Ohm) doesn't matter if both lv and thresh. fanout is plugged, does matter if LV power plugged into box before pair 8 was fixed 5.60M no LV power 32k no LV power after pair 8 fixed 200 with LV power MB2 Top Beam Right Total=2kOhm because skipped on connector, pair 25 is used on this one 1 open 1bar 25 14 18 2.8M 2bar 24* was open 4 18 46 10k 8k 3bar 23 5 84 43 540k 570k 4bar 22 13* 56 42 open, push on tight, very flakey 14* 76 30 open, push on tight, flakey 18 16 21 26k 25k 17bar 9 20 13 29 16k 24k 19bar 7 23 94 32 5k 4K 22bar 4 25 39 27 577k 500k 24bar 2 open 25bar 1 * = fixed Connected HV to Lecroy 1440 1443PF (positive) Modules (16 channel) slot HV cable ----- ------------- 6 1-16 5 17-32 4 33-48 3 49-64 2 65-80 1 80-96 Possible leak at fitting of valve on output of flow controller (snoop-it) Gas system normal operation with CO2 o about 10psi on gauge on supply o both highflow and lowflow valves open o both rotameters (high and veryhigh) closed o flowcontroller actively metering gas o pressure relief valve at top of detector set at .25 inches (so it is possible to waste gas if flow controller is set too high and top valve is venting) o pressure gauge at bottom of detector reads ~ 0.5 inches o note that the scale on the top/bottom gaugesare different
February 11-12, 2004
miniboxes now have 2x 3Ampr fuses (one fuse for 12 channels)
Bob Jones put the LV fanout connector back in upside down, redid myself
MB1 by itself
-5.66 Volts no load
-4.60 Volts with all 24 bases
~1.5Amps according to crude gauge on power supply
MB2 by istelf
-5.68 Volts no load
-4.60 Volts with all 24 bases
~1.5Amps according to power supply
-4.6 Volts with all 48 bases
~3.5 Amps according to power supply
-5.63V at power supply
MB3 -4.67V on bases
MB4 -4.64 V on bases
Current gauge does not appear to work on this supply
Raised both power supplies to -5.68V (max it will go)
Wrote hvc program for controlling hv, in use by Mike Longo and
March 8-10, 2004 Fixed branch highway to CKOV crates Rack 6, VME crate contains beamppc and camac controller connects to crate 1 Rack 3 -> crate 2 Rack 6 -> crate 3 rack 12 -> crate 4 rack 13 CKOV crates are 3&4 Branch highway from rack 6 to rack 2 is made from two 25 foot cables with 1foot union the union is in cable tray opposite rack 7 Removed unioin o one of the pins was not fully inserted, disassembled connector and repaired o screws on face of connector just barely touch those on the other connector replaced screws on one end of union with lower profile, now does not touch Three low bits were always zero. After fiddling with union ok, but, address high bit (=16) not addressable push on cable going into crate 4, now everything ok Testing the branch highway Logon to e907daq and then ssh to beamppc (use your own account or daq if you don't have one) source /usr/local/mipp/setup-mipp-online.csh initCAMAC 1 [crate] execCAMAC 1 [crate] [slot] 17 1 [value] # write value to slot in crate execCAMAC 1 [crate] [slot] 1 1 # read value from slot in crate Matt, (Grad. student or postdoc?) from Colorado working on TOF Fixed 1443PF HV in 1440. Problem was that HV readback was only tens of volts at HV settings of few hundred, it turns out it was caused by positive current limit (read with RL, set with LI 255) HV power is supplied by 1442 modules plugged into mainframe. There are two bays for these modules on the left side of the unit. The left-most module was missing, there was a face plate with LBL stickers, but no unit! Got two 1442 modules from prep, installed one and labelled the other as spare. Also returned one 1442 module to prep that Carl said was bad. Note that the ribbon connector is keyed, but you plug it in the other way. If ribbon cable is up-side down, some of the LEDs on the 1445 will flash slowly instead of being steady on. The mainframe modules tilt forward. Unscrew only the top and tilt down. There is a spring loaded stop that holds the unit open. Very convenient. Hank the prep HV tech said that the left-most 1442 powers slots 0-7, with both positive and negative. The manual says otherwise, I did not verify this. Analog signal check on channel 84 At the base, need digital scope to see signal (can use analog scope but can only measure the rate by looking at trigger light, won't be able to see the trace) 50 Ohm DC, -30mV trigger at 1800V see 10nsec wide pulse at about a hertz At the end of the analog cable, the signal is a bit fatter (from the delay line), lower threshold to -25mV On output of delay boards signal is really small and very long (tested without ribbon plugged in, lemo output is on solder side, note that the ribbon and lemo are just shorted, remember that our boards are different than the TOF boards) go to .1microsec per division and trigger at -5mV See lots of ringing in signal now, impedances are not matched. ADC counts should be the same. Threshold DAC BiRa 5408 on beamppc run ~lange6/Helpers/bin/test5408, e.g. test5408 1 4 16 0 -.5 read back should be same as set voltage Tested output on lemo going into minibox, within .001 of DAC setting. Tested output at subminix cable going to base. Note that with Low voltage on, there was -.4 volts on the threshold line even with the threshold set to zero. thresh lv off lv on 0.00 0 -.401 -0.50 -.483 -.590 -0.75 -.724 -1.00 -.960 -1.042 -1.25 -1.195 -1.50 -1.423 -1.490 -1.75 -1.641 -2.00 -1.735* -1.935 * going down down (towards zero) .001 every 15 seconds Card locations Crate 4 2,3,4 ADC FERA 4300 6&7, 9&10, 12&13 Disc 4415A, TDC 3377 14 Disc 4413 (lemo in, rats) 15 Scaler 2551 16 DAC 5408 17 NIM to ECL RFD ? 18&19 Gate Delay 2323 Crate 3 2,3,4 ADC FERA 4300 6&7, 9&10, 12&13 Disc 4415A, TDC 3377
Notes from March 15-19, 2004 trip Rearranged crate 4 14 NIM to ECL 15&16 Gate Delay 2323 17 DAC 5408 18 Scaler 4434 19 Scaler 2551 20 Disc 4415A 21 Pulser Must initialize crates for 2551 scaler to work!, i.e. initCAMAC 1 4 Crate 4 top address bit dropped again. Touched branch highway cables in rack 6 seemed to fix it Checked digital output on scope Trigger on analog output (directly from base) 50 Ohm DC, 50mV, 10nsec, -15 mV threshold Short (8foot) ribbon cable from spider box to ECL-NIM modules in NIM carte Look at NIM output of ECL-NIM on digital scope Channel 84 (bottom) MB3, Spider box 4, channel 12 HV slot 1 channel 3 (counting from 0) digital output has -300 mV DC offset (at 50Ohm) (NB through ECL-NIM module) digital signal shoots back to zero volts and then down to -2V. jagged and triangular with a width of about 150 nsec Digital signal after long ribbon cable (to rack 13) and then to NIM-ECL -300mV offset, big positive pulse (-300 to 0 mv) width ~40 nsec followed by big negative pulse (-300 to -500 mV) width ~40 nsec, but 150 nsec after first pulse Digital lines must be inverted! Invert ribbon cable (now find Channel 84 on channel 6), DC offset is gone, nice negative pulse -0.5 volts, primary width 50nsec, but multiple trailing pulses of diminishing magnitude, typically see two smaller pulses after main one. Analog signal on channel 84 sometimes disappears. Touching subminex connectors on the base seems to solve this problem. CAMAC has intermittent problem with highest address bit. Touching rack 6 branch highway appears to fix it. Digitial output does not appear unless threshold set to -1.8 (For Channel 84 this is MB3 SB4). With new frankenstein cables here is how to get digital signal Both + - should have 56Ohm pulldown resistor to -2V (Note that True+ is white base output labelled -, True - is blue base output labelled +) Make sure threshold is set to -1.8V and LV and HV are on. To see signal on scope, pull out + or - signal (with powered pulldowns) and couple scope with AC 1Mohm. Will see 1.5V by 40 nsec wide pulse (positive going for + and negavite going for -, note that there is a DC offset that keeps the signals entirely negative) Note that there is a smaller (~10 times smaller) signal that is 200nsec after the real signal. This is probably a reflectiioin somewhere along the cable. This is probably not due to the ribbon to lemo conversion on the scope because if you put the ribbon directly into the discriminator you can lower the threshold enough to trigger on both the big and little signals. Set all Discriminators (4415A) to minimum (100nsec) and 300mV (on test port). If truly differential this is equivalent to 150mV. The frankenstein cable should be male with female close by to accept terminating plug, then female on the other end. Leave space between the male and female end, because the female end must duck unto cables near the discriminator input. LED bench tests Single 325 Ohm resistor in series with n LEDs in parallel Green LED is visible at 1.80 Volts, supposed to be 20 mA at 2.2VF Yellow is 10 mA at 2.2VF testing green LED Vbias Vresistor nLeD 1 2 5 V mV mV mV 1.80 41 58 77 1.90 101 125 148 2.00 179 203 236 2.10 259 286 321 World's greatest pulser: 1340 Pulse generator (from Prep) has variable output 0-10V as well as pulse width and frequency. Provided that pulse generator and NIM linear fanout can drive LEDs Reorganized DAQ crates again Slot Crate 4 Crate 4 1&2 2323A Gate Delay Empty 3-5 4300 ADC Same 6 RFD04 NIMto ECL Same 7&8 4415A Disc/3377TDC Same 9&10 " " 11&12 " " 13&14 4433 Scaler Same 15 Empty 16 5408 DAC 17&18 Empty 19 2551 Scaler 20 4416B Disc Note that third ADC and TDC in each crate should be terminated (for gate chain). Modified jumpers on 2323A so that ECL output is Delayed (jumper inside, see manual) All Modules (ADC,TDC,Sclaers) are working and can be readout. 4415A discriminator works on base digital output with -2V through 56Ohm pulldowns. Split analog signal from base, one half to 4416 discriminator, other half to ADC. See charge on ADC before and after delay boards from South Carolina!
Notes from March 30-April3, 2004 trip (Lange) o Hardware Found that I could not control the dac. Did not appear to be a problem with camac bits being dropped. Put in a 'spare', which worked, but does not have the right voltage range (need -10-10, spares are 0-10). Took both the working and broken modules to prep for repair. Say that early next week will be done. Will ask Doug to get them. Plugged in the digital output with swap cable directly into one tdc. See lots of hits (though threshold is not connected...) This might indicate that tdc would work for us without kludge cable Talked to Mike B. about discriminator inputs...Didn't have a solution (expected it to work, save our strange ecl signals) Got from Mike 24 10 pin 56 ohm sips. Basically cleaned them out... Also found some cable and a few (but not enough) connectors. Did not make it to the stockroom to see if they have the male connectors or not. Its really hot near our crates - in fact they are hot to the touch. Should ask about getting a fan in that area Got a trigger line .. not labeled yet. Found that we were not seeing the beam trigger pulse. TOF also was not. Traced by Andre/Pierrick to be a very short pulse (5ns). Now fixed. Both 2323a modules ECL for channel 0 is the delay. For the 2323a in crate 3, ECL for channel 1 is not delayed. For the 2323a in crate 4, it is. I'm not sure how the trigger timing is working. Given that the minimum delay is 50 ns we may need to use the non-delayed output of the 2323a. Kludge cables are made. 1 is connected to 1 4 12 tdc the other 5 are in the plastic bag in the box in the counting house. o Where to find stuff in daq-land nominal ckov config file ::/ppcboot/common/usr/local/mipp/online/config/ckov_config.xml log file from current run ::/ppcboot/common/usr/local/mipp/log/ckov.log log file from previous runs ::/ppcboot/common/usr/local/mipp/log/archive/ckov.log.*.gz o Software: --> Got CkovReadout executable working. Defined data format for the CKOV: Code lives in Detector/Ckov/CkovDetector.cxx, needs to be consistent with MippIo/MippIoCkovBlock.cxx also. // 0xADC // number of ADC words read out // wordcount*0x800 + crate*0x100 + slot of the ADC // ADC word 1 // ... // ADC word N // ........... for other ADC's // wordcount*0x800 + crate*0x100 + slot of the ADC // ADC word 1 // ... // ADC word N // 0x2DC0 // number of TDC words read out // wordcount*0x800 + crate*0x100 + slot of the TDC // TDC word 1 // ... // TDC word N // ........... for other TDC's // wordcount*0x800 + crate*0x100 + slot of the TDC // TDC word 1 // ... // TDC word N --> Learned how to commit code to production area --> for some packages there are two areas /ppcboot/common/usr/local/mipp/online (powerpc code) /usr/local/mipp/online (e907daq code) --> in my /ppcboot/common/home/lange6/prod and /home/lange6/online/ areas I have set up compiling areas to allow for easily updating production code (ie, make install is a dangerous command in these areas) --> Added code to RawData, MippIo, Raw2Root for parsing the raw data format of the ckov. Most of it is taked from the tof code, which is nominally working. --> setup environment: cd /home/lange6/mippSoft/040402 setenv code_907 /afs/fnal.gov/files/code/e907 source $code_907/mippsoft/bin/setup_linux.csh --> Command for Raw2Root (convert raw format data into a root file) bin/Linux2-GCC/raw2root /data/1/mipp00004938.0000.ra --> Command to run over raw root file bin/Linux2-GCC/anamipp -i raw2root.root -x JobControl/xml/test/job.xml -o test1.root -n 100 --> Wrote a simple module to make a th1f of my adc counts in /home/lange6/mippSoft/040402/JobControl/test/CkovTest.cc Need to find a home for ckov analysis code. o Co2 --> Terry gave me the quick tour of the gas computer. Shows a nominal rate of 4 ft^3/day. Need to check the variation of atm. pressure. Plot of co2 use vs. time shows significant variation with time. Rate of usage is 3x projected. May be an issue given c4f10 cost. Did not manage to re-leakcheck the ckov. Should ask if we know that mass-flow usage is consistent with bottle usage. (may not be known...) o HV --> Problems with Andrew's HV program. Trips channels off that are more than 10V from nominal. However, a program setting of 1800 volts results in a reading of 1807-1812, so this is a problem for us. Turning off the hv_monitor program is a temp. fix. As of 040403 1am, Andrew has loosened the limits for the ckov channels. --> Our HV is still off by default. Can be loaded via ckov.hv file. o Data --> a few 'long' runs with ckov at 1800V and beam. Nominally run 4938 is a 'good' run with the ckov running. The delay wa 50 ns, delayed width is 300ns. Have not checked that this is 'ok' or not. --> Run 4971 is 50K of pulser triggers. Trigger prescales were 500 for pulse generator prescale (at the top) 20 for pulse gen 1 prescale (at the bottom) 20 for pulse gen 2 prescale (at the bottom) this gave 10s of Hz, which is about what the daq seems to handle. --> Run 4972 is taken just after 4972 with the ckov HV on is with beam. data for each pmt plots email
Notes from April 7-9, 2004 trip (Wright) Discriminators in slot 17&18 of Crate 4 go with TDCs in Crate 3 In Co2, the pion threshold is about 5 GeV, so in a 40 GeV beam, at least the primary particles should radiate, except for protons. LV was off for Ckov top, switched on at 7:55pm 4-7-04 (bottom was already on) config file for ckov daq (controls gate delay and width is here) /ppcboot/common/usr/local/mipp/online/config/ckov_config.xml Gatelen = gate delay (ns) DelayWid = gate width (ns) Run Control FAQ trigger prescalar is n+1, i.e. 0 = scale of 1, i.e., none 1 = scale of 2 i.e., every other trigger beam trigger fails to work if nothing running on richppc (richppc is then never ready) Ckov DAQ takes microseconds 10^3.1 e907 daq setup required to run david's analysis programs (anamipp...) cd /home/lange6/mippSoft/040402 setenv code_907 /afs/fnal.gov/files/code/e907 source $code_907/mippsoft/bin/setup_linux.csh /ppcboot/common/usr/local/mipp/online/config/ckov_config.xml took beam data with various gate delays width 50ns run delay events 5068 0 108? 5069 50 122 5070 100 118 5071 150 123 5072 200 22 5075 200 5076 200 5077 250 5079 300 5080 350 5081 400 "fixed" DAC BiRa 5408 from prep don't work at all neither addressable from CAMAC (red light should come on) one module puts out 0V by default, the other puts out -2V by default left -2V DAC installed (Crate 4, slot 15, to drive thresholds) Fixed trigger by magic hands touching logic module Lecroy 2365, slot 9 of branch 1, crate 1 (Rack 3) This module doesn't like being all the way in tight. use testLecroy3265 1 1 9 If you get error syncing it isn't working. hv crates not responding after reboot of e907daq minicom couldn't contact either reset serial port (button on front) there is no power switch minicom can now access hv but not hvc or hv_control restore proper settings on the device files: # stty -F /dev/ttyQ01e6 1:0:8bd:0:3:1c:7f:15:4:5:1:0:11:13:1a:0:12:f:17:16:0:0:0:0:0:0:0:0:0:0:0:0:0:0:0:0 # stty -F /dev/ttyQ01ec ... now it works returned -10 to 10V Bira 5408 DAC to prep, asked them to check ability to read from CAMAC
Notes from April 14-17, 2004 trip (Lange) o For next trip (Dougs) Items to check during access: o spider box to tdc map o output of 5408 module. Still -2.0? o Are all threshold lemos plugged in at detector end? o Lam problems in crate 3: o 1000 event pulser run with ~20Hz rate should show that some times we don't get all lams. You can see this in the lower right plot on rc-gui if there are 4 lumps in the ckov readout time. If not, check the log file as well. o Swap cables for adcs. Does problem with lams follow cable? o Swap 2323as. Does problem with lams follow 2323? (I doubt it) o Take the terminated adc in slot 5 out of the chain (change the config file and unplug the ribbon cable for the gate. If this solves in then unterminate that adc. I don't quite remember how, likely some resistors in the upper part of the board (inside..) There is no terminated adc in crate 4. o Double check that hex bolts on side of ckov are tight. Look for gas leaks.. o Items to follow up on without access o 5408 status at prep? Do we need a different solution? Eg, Power supply in counting house? Progress: Fixed readout and reconstruction code for the TDCs. Any data before run 5222 is junk for the TDC readout. Had truncated the # of words per tdc at 32. The data is there, but it will take some kludging to process the raw bits. Starting from 5222 the data can be read out in a sensible way. Found that the 5408 was reading ~0. Power cycling the crate returned it to -2V as Doug saw originally.. Verified that I can not talk to it via camac commands. Prep has not verified the functionality of the other 5408. Argh. o Further fixes to adc and tdc readout code. Added map of mirror #, adc channel, spider box, tdc to database and now use map in Raw2Root code. Map can be found at http://cdcvs0.fnal.gov/cgi-bin/public-cvs/cvsweb-public.cgi/e907/mipp/offline/CkovReco/maps/ note that it has not been checked against the cabling scheme. o Have started a CkovReco package for my histo module. Andre updated several things just after I stopped working, so need to resync to stay up to date However, the code I have in ~lange6/mippSoft/040402 is working. The recipe is unchanged with one exception, run anamipp with a different xml file as input raw2root -o
Notes from April 20-23, 2004 trip (Wright) Visited prep, BiRa 5408 repair was not in the priority que as it should, talked with Tim, will be fixed tomorrow Confirmed that threshold lemos are plugged into all Miniboxes Burndy 30P NE, need female connector for LED box Got Lange's raw2root to work for me, but have problem with database running anamip Got 5408 working in slot 21 of Crate 4. Slots 15-17 do not work with the two 5408's recently returned from prep. This appears to be a problem with the crate. Can definitley feel a difference when plugging in the 5408 modules. Much more solid fit in slot 21. 5408 CAMAC commands Set voltage: F16 Channel(0-7) Value(0-4095) 0=-10V, 2048=0V, 4095=10V Read voltage: F0 Channel(0-7) LED bias voltages (measured with DVM in diode mode) note long lead is +, short lead is gnd green 1.770 yellow 1.705 red 1.529 ckov led 1.75 - 1.76 Burndy connector lables pins A-Za-z, skips I,O,Q + -(gnd) +-(gnd) AB NP CD RS EF TU GH VW JK XY LM Za Installed linear fanout in NIM crate immediately downstream of RR13 (labelled module CKOV LED) Set threshold on 5408 to -2V (1,2,3,4) 10am 4-23-04 /ppcboot/common/usr/local/mipp/online/config/ckov_config.xml changed DAC from slot 16 to 21
Notes from April 30-May 4, 2004 trip (Lange) Visited Bob Jones: led box is in line behind tpc and rich work. Also, Bob had a question on the design, which I answered incorrectly. Doug followed up to fix the misinformation and answer other questions. Exchanged non-functioning 4300 with prep. Talked to Terry about the c4f10 status. Aparently company send c3f8 and is figuring out how to remedy the situation. No estimate for time yet. Not yet urgent from our standpoint. Checked ckov log files. See that no longer getting any lams from crate 3. run 5440 seems ok, 5481 is not (problem on apr 21 or 22) The daq is back as of 4pm Saturday. Problems with missing lams gone - at least with pulser trigger. Found that the SP2 kludge cable was a) plugged into the tdc upside down, and b) not flipped. Result was that the channels connected to sp2 had no tdc hits (with one exception). As of run 5619, this is fixed. Raised ckov hv to 2100 volts for tonights running. Will need to email Andrew about increasing the overvoltage point. Andrew since reminded me that this is controllable in setup files in ~daq/hv_control/voltage_limits.hv. Ckov limits now 30 (over) and 50 (critical). TDC default resolution is 0.5ns (set in SetCommonStop method), so peak at 400 counts = digital signals arrive 200 ns before trigger (+2323). Hum - this might indicate that even 50ns of delay is too much for adcs. Need to look at adc data (crate 3 has no delay in adc gate, crate 4 has 50 ns delay). PH changed our trigger to come from a different source (next to Rosie) - looking at data with this trigger, we had no signals in our detector. Access today, I checked the relative timing of these signals. Found that the 'new' trigger comes 500ns (!!!) later than the 'old' one. That would explain it.
Notes from June 1-3, 2004 trip (Wright)
Picked up LED fanout box, tested ok
but discovered polarity is wrong for -2V supply!
Linear fanouts only work negative.
Using positive output of 1340 pulse generator to
drive all bottom LEDS (put test LED in connector for top
to see light output) can see increased output on analog
output of PMT 81 at 1800V.
Pulse settings delay ~ 1.5microsec, width = 1microsec
amplitude is about 2V,
NOTE: amplitude has two knobs (red and white lined)
red controls the positive output, white controls the negative output!
Performed surgery on pulser box, swapped purple and green wires
going to/from resistors, now can use -2V, but now switch won't work
since the Vin lines are now all tied together.
1700V PMT 81 (bottom, way in the corner).
NIM Linear Fanout will now drive LEDs, in order to trigger on LED,
must use output of second fanout of quad. If you try to use same
fanout, can not trigger on the scope (everything is set to 50ohm input
If you are driving only bottom, then get light with input voltage
starting at -1.8V, the maximum output seems to be at -5V (remember
fanout will limit the voltage). If you are driving both top and
bottom, no light output below -2.5V.
Note Pulse width must be greater than 1 microseconds or you get no
analog output. Can run frequency as low as you want (even 1Hz or less).
First light arrival is 1.5 microseconds after start of pulse.
Victory! bottom 1800 V
3000 ns gate 0 delay
pulser 100hz 2.5 micro width, amplitude -2.25 v on scope (3 on dial) inside
-2.0V on scope in counting ouse (-2.25V in MC7) (3.5 on dial) counting house
Found DAQ problem in crate 3. ADCs would not produce LAM. Crate 4 ok.
Turns out problem is caused by too many trigger modules, crate can't
put out enough power.
LED box design (pdf)
Notes from July 13-16, 2004 trip (Wright) Crate 4 in strange state. Crate power was on but crate controller and branch highway terminator (with the lights) were completely out (no lights on). Cycling the power restored to normal. DCH Trigger modules were still stuffed into crate 3 of our system. Andrew Norman and Ken Nelson are working on installing new crate immediately upstream (small black minirack). After brief discussion with Andrew, I left all trigger stuff cabled, but pulled modules so that not drawing power. DCH using our NIM to ECL Camac module. We use it for timing TDCs. They replaced the cable so that it drives both our TDCs and their MALUs. Could not find original, cut DCH tail (there is enough room for them to crimp on new connector). Tested ADC modules (4300) in Crate 3 OK used: i; w 3 3 255 w 3 4 255 w 3 5 255 Plugged in Disc. modules (4415A) in Crate3 (modules were in place, but not plugged in.) Tested ADC modules (4300) in Crate 3 OK Tested ADCs using Gate/Delay OK used: pulse # generates 10hz using 2323 module 4300 1 3 3 10 4300 1 3 4 10 4300 1 3 5 10 4300 1 4 3 10 4300 1 4 4 10 4300 1 4 5 10 Someone removed the power strip next to crates 3 and 4. Cable tied new power strip to rack, replaced long extension cord with small one, cable tied in place. Analog 350MHz scope was bad, weird artifacts on traces. Labelled scope as bad, replaced with 300MHz analog scope. Rewired NIM linear fan-out that drives the LEDs to make more obvious and provide more output lines. Output of pulsers (one in worm and one in counting house) now go to the input of the bottom fan-out. The output of the bottom fan-out now drives the top two fan-outs. The fanouts need terminators on three of the four outputs (see manual) and at least two of the outputs must be terminated. Put two terminators on all channels. Andrew's hv_control and hv_monitor now run on e907hv.fnal.gov in counting house. Each program runs in a separate xterm, can't tell them apart from the icons, so I copied my 'name' script to /homehv/daq/bin and named each xterm (monitor and control). Found bug in hv_monitor, can't go back to CKOV display because there is an empty card between last card and next system. Asked Andrew to fix this. Also requested he implement the HV channel to PMT map. Serial ports for HV controller are now on e907hv and not e907daq. Must run hv_control on e907hv. Andrew et al removed trigger cards from crate 3, now all hell has broken loose. Neither crate will work. Crate 3 bottom branch highway cable was a little loose, now crate 3 works. Crate 4 has a variety of bizarre problems. Access light (lights up when card is being addressed) is latched on for the 5408 and one of the TDCs. This is an error, should only happen when DAQ addresses the card. The ADCs no longer respond correctly. If you write a value, the read back is incorrect. For example, w 4 3 255, reads back 169. The 2323A also not working properly. Pushing "CHAN" toggle causes display to go "crazy". Trying to pulse the 2323A with script 'pulse' causes the output to latch on at high rate. Problem goes away if you plug module into another crate. If I unplug all modules on crate 4 except crate controller, terminator and one ADC, the ADC works. If add another ADC or any other card, the ADC no longer work. Can read them, but get wrong value. If 5408 is only card in crate (not counting controllers) the access light is out, add any additional card and the access light on the 5408 latches on. Note that crate 3 nominally works, but reading back ADCs gives the wrong answer about 1 in 5 tries. Crate 3 card 3 seems to be the worst. So, replaced crate 4 with new CAMAC crate. Talked to Mike Benke while at prep, Mike suggested that high temp and high humidity could be cause of problems. With new 2323A (from prep) and original three ADC modules and single 5408. Crate 4 is working so far. But... crate 3 has problems now! ADC modules give incorrect readbacks about once in 5 tries. 2323A goes crazy as described in crate 4 story above. Spare 2323A works fine in crate 4, but does not pulse correctly (via pulse script) in crate 3. So, removed first ADC module from crate 3 and put in crate 4. This module works fine. Unplugged all modules from crate 3. Single ADC in slot 4 showed periodic readback problems. Wrote script "darn" to repeatedly write to a single card. Trying script now shows that card 4 is fine! Try by hand again and it is still fine. Now added card in slot 5, ok. Now put back card into slot 3, still fine. Hurray! Crate 3, put 2323A back in. Now it is fine, not crazy and pulses at correct rate. It is noticable colder now in the worm (it is 6pm) than it was earlier today, perhaps this is why things are working again. Wrote script 'testadc' to run pulser and readout all ADCs on both crates. Both crates are ok. Note that only 2323A, ADCs and 5408 are plugged in. Asked David to change test4300 so that min, max, avg outputs are in columns. Found bug in test4300, always got first trigger even without real trigger, probably due to test command in program, David fixed this. 4300all [trig] [file] command will take triggers from all adcs and put results in file, each column is output of one channel. Reinstalled LED pulser unit in counting house. Somebody removed it to install a bunch of scalers. Cabled both (crate 3 and 4) 2323A so that can get triggers from LED pulser and from official trigger, inputs will drive both 2323s. Tested both LED pulsers (in worm and in counting house) to drive 2323, ok. Crate 3, 2323 trigger threshold set (using trim pot on front panel) Chan A threshold -0.300V B -0.303V Trigger toggles should be set to negative(-) slope How to program the 2323A: 2323 1 crate slot chan delay gate 0 0 chan=0,1 = A,B NIM output triggers at same time as START and has width is equal to progammed delay DLY output is delayed by programmed amount with width 10, 30, 100, 300ns ECL output is either NIM or DLY, depending on internal jumpers Our modules should be set to NIM. Verified this for crate 3. The reason for this (according to David) is that the minimum DLY is about 50ns and the experiment trigger comes so late, that we can't afford this much delay! Found bug in test2323, always resets chanA and chanB, so can't program them separately. Fixed by David. Recables gates so that LED goes to chanA on crate4, DLY output triggers chanB which is daisyed to chanB crate 3. The experiment trigger is also in crate 3 chanB. So can now program crate 4 to generate pulse with arbitrary delay (chanA) with arbitrary width (chanB). Wrote scripte gate to set crate 4 chanA and chanB. Verified on scope that this works. Wrote test_dac script to test 5408 (slot 20 in crate 4). Cables threshold lines. Test ok. Installed NIM to ECL (to drive TDCs) and TDCs is crates 3 and 4: NIMtoECL slot 7 TDCs slots 9,11,13 Wrote test_tdc to read them out. All ok (both crates). Note that Discrimantors and scalars are not plugged in. Cabled ADCs and TDCs is both crates. Wrote test_all script to test ADCs, TDCs, and DAC. All ok! Updated ckov_config.xml (/ppcboot/common/usr/local/mipp/online/config) changed TDC slots to 9,11,13 and DAC to slot 20. 11:40am Although it is still cool on an absolute scale, it is starting to get noticeable warmer by our racks. 1:00pm A bit warmer. Getting blasted with external trigger, this interferes with test_all pulser. Unplug external trigger, everything ok. Can tell if getting blasted because ADCs readback very quickly before hanging up. Andrew is running a pulser on the experimental trigger. This is automatically fanned out to everyone! I asked for trigger redesign, so that can program (or disable) each detector trigger via computer. Trigger guys insist on getting some data today. Ended access around 2pm? and waiting around MCC to generate beam. They have lots of problems, so finally found someone to do controlled access with me. Timing tests. Put digital scope by LED pulser. With three equal length cables, connect scope (chan1) to common input to 2323 Crate 4 ChanA, (chan2) to NIM output, (chan3) and to analog output of PMT 81 (must unplug from delay board). Set pulser to 30 Hz. Using chan1, set pulser output voltage to 3V (beware, analog scope was not calibrated properly and read only 2V, I believe the digital scope because it correctly shows -1V on NIM output). Also using chan1 set pulser width to 1.5 microseconds. Set PMT to 2000V. Will now see lots of PMT pulses (chan3) begining about 1.25 microseconds after beginning of chan1. The main pulse train is about 1-1.5 microsec. long. There are large single peaks at around 1microsec. Can now use 'gate' command to move ADG gate. Note that the delay boards introduce an additional delay of 300ns. Can look at signal before and after delay. Put tee in PMT 81 on input to delay board. Note that there is an additional lemo socket on the outside of the delay board, this is tied to the delay output. Connect (chan3) of scope to the tee and (chan4) to the back of the delay board. Both should be 50 Ohm DC coupled. Chan3 set to 10mv and Chan 4 to 5mv. With this setup, can trigger on big spikes on chan3 and can measure delay on chan4, verified that delay is 300 nsec. Pedestal scans -------------- With HV off, check pedestals for various gate widths. Some pedestals increase and some decrease! as width goes from 100 to 3000 nsec. HV ADC slot crate slot 1 4 5 2 4 4 3 4 3 4 3 5 5 3 4 6 3 3 Pulser in counting house. Amplitude set to 6 (on the dial). Duration set to 1.5 microsec (on the dial). Frequency set to 400 hz (on the dial). Trigger for 100 events. Did delay scan for various widths. Good setting is gate 900 1200 (delay width). Friday, July 16, 11am. Very warm in worm. Closed all night and had beam. Crate 3 and 4 now have severe problems: test_all fails because of crate3. 2323 showing pulser failure mode. The display at top is flashing (wrong behavior) and the output frequently gets stuck on at high rate (should only be 10hz). Note that 2323 and ADCs are extremely hot (feel at top of rack). Other modules do not feel very warm. ADC in slot 3 on both crate 3 & 4 shows read-back problem (using 'darn' do not get full counts, i.e. 255). The other ADCs are ok, a ha, perhaps this is the cause of the failure! Moved ADCs down one slot, to leave gap between ADC and 2323. Note I moved 5 to 6 and 3 to 5, leaving 4 as it is. Problem with ADC readout moved with the card, i.e. now readout problem is with slot 5. Leave this for now and move on. 2323 is not bad, must be crate 3 too hot, because moving the module to crate 4, it works fine, moving it back to crate 3, it has problems. This must be the bad connection to crate bus that Mike Benke talked about. So, replaced crate 3 with new crate BiRa 6700. 2323 now works fine. ADC modules in new crate have read-back problem, i.e. writing and reading register does not result in same number. It looks all but the lowest two or three bits not working. Tried four different ADC modules in many different slots. They work in crate 4, but not crate 3. It must be the crate controller! Replaced Jorway 70A crontoller with spare, and presto the ADCs work! It looks like the old crate controller had problems with the high bits on the dataway. Put in all ADCs and TDCs in crate 3. All modules read back correctly. All modules getting proper triggers from 2323. Did pedestal measurements on ADCs. Note that two modules giving abnormally low pedestals (0-2 instead of 20-40). Problem appears in middle ADC of crate 4 and last ADC of crate 3. Turn on HV and do LED pulser runs. Can see signal on all ADC modules, except the two above that had pedestal problems! Always reading 0-2 counts. Replaced modules with spares (note that I had to move the terminating resistors from the existing module to the spare). Now getting reasonable pedestals and seeing signal on all ADCs. It is still hot in the worm, but everything is working! Edited ckov_config.xml to match new slot scheme. Set thresholds on both 2323 to -0.2V. Note that ChanA of Crate 4 must have terminator on unused input, primary input comes from linear fanout of led pulser signal. Both 2323 ChanB ECL output should be jumpered to DLY. Verified that pulsers inside worm and in counting house work (getting triggers). Time to catch my flight. Let's hope nobody touches this after I go.
Notes from Sep, 2004 (Wright from LLNL) 9-14-04 DAQ people unnecessarily changed CKOV crate addresses from 3,4 to 5,6. This causes CKOV DAQ and all test scripts to break (total of about 20 scripts and files between Lange, Wright, and official DAQ). Prevention: DAQ people should have used 5,6 for new non-CKOV crates, thus CKOV would not need any changes. DAQ people unnecessarily removed execCAMAC command, replaced it with two different commands (exec16, exec24). This breaks subset of CKOV test scripts. Also functionality of commands is now different. Could not use trial-and-error to figure out which command to use since entire racks were left off after power failure (no crates responding). According to Andre (via phone) exec16 is equivalent to old execCAMAC. Prevention: Should have replaced old execCAMAC version with link or some other backward compatible fix (to use exec16) and added execCAMAC24, thus CKOV would not need any changes. Power to racks 12 and 13 was not restored after shutdown. Consequently can not read out CKOV crates. Had to debug system (since software changed), once discoverrf that system really isn't there, called control room and ask someone to turn on power. Prevention: DAQ and/or run manager should have had plan to restore system racks to state prior to power shutdown. Would have saved CKOV debugging time and random onsite person to fix problem. 'test_dac' now works as advertised. Note that this is in crate 6 (formerly known as 4) > test_dac setting to -0.5 volts Read back -0.50061 Read back -0.50061 Read back -0.50061 Read back -0.50061 setting to 0 volts Read back -0.002442 Read back -0.002442 Read back -0.002442 Read back -0.002442 Wrote test_there script to write and read back data ADC and TDC registers. ADC and TDCs respond: > test_there writing 255 ADCs crate 5, slots 4,6,8 Data = 255 = 0xff, Q = 1, X = 1, timeout = 0 Data = 255 = 0xff, Q = 1, X = 1, timeout = 0 Data = 255 = 0xff, Q = 1, X = 1, timeout = 0 ADCs crate 6, slots 4,6,8 Data = 255 = 0xff, Q = 1, X = 1, timeout = 0 Data = 255 = 0xff, Q = 1, X = 1, timeout = 0 Data = 255 = 0xff, Q = 1, X = 1, timeout = 0 TDCs crate 5, slots 12,14,16 Data = 255 = 0xff, Q = 0, X = 1, timeout = 0 Data = 255 = 0xff, Q = 0, X = 1, timeout = 0 Data = 255 = 0xff, Q = 0, X = 1, timeout = 0 TDCs crate 6, slots 12,14,16 Data = 255 = 0xff, Q = 0, X = 1, timeout = 0 Data = 255 = 0xff, Q = 0, X = 1, timeout = 0 Data = 255 = 0xff, Q = 0, X = 1, timeout = 0
Notes from September 21-22, 2004 (Wright) Debug crate 6 because not working Found bad 70A, replaced with good from Prep (two spares from hall were bad, returned these to prep.) Test sequence to see that modules are there test_dac test_there 255 test_there 111 test_there 0 OK all modules present and accounted for. But, ADC slot 6, crate 6, is flakey, read back test doesn't match data 111 do get triggers and reasonable pedestals Replaced with spare marked possibly bad, read back test ok, but pedestal values now nearly all zero. Put original ADC module back in. Keep an eye on this. Edited pulse, pulse_camac to work with new crate IDs. test_adc 6 ok, get triggers and see pedestals test_adc 5 fails, get triggers without pulser! test_tdc 6 ok, get triggers and see data test_tdc 5 fails, get triggers without pulers, all null data! crate 5 controller was bad too! Probably when our colleagues changed our crate id's from 3,4 to 5,6 they just took our good controllers and gave us crap ones! Trigger input to 2323 crate 5 needed 50ohm terminator. Exp. trigger was on, disabled output register from DAQ (required run start to actually download). Now all triggers (pulser, LED, experiment) are working again! Rewrote scripts test_dac, test_adc, test_tdc, test_there, test_all to add features. Wrote test_led script. Use with LED pulser to see signals on ADCs. Results, using LED pulser in worm: delay= 900 width= 1200 ns set HV to 1900V (N.B. wait at least minute for bases to warm up) set pulser: frequency = 400 Hz duration = 2 microseconds amplitude = 6V (white, 0V red) if in counting house 5V (white, 0V red) if in worm offset = 0v (white and red) negative pulse output to LEDs Check that ADCs get triggers: You should see six lines that end with 'Average' 110 59 357 109 58 32 1210 96 28 52 605 54 219 52 737 52 Average 67 35 217 82 1019 50 2047 69 273 69 1135 70 480 88 370 24 Average 2047 49 293 135 92 62 2047 180 148 110 150 106 68 85 0 89 Average 109 56 4 31 26 45 180 46 111 51 173 43 202 26 562 5 Average 1623 28 649 67 786 83 74 851 72 59 334 75 355 71 559 26 Average 32 42 136 41 84 88 116 0 68 115 180 26 23 602 37 27 Average Tested spare 4300 ADC boards (3 of them) with LED pulser, none of them worked, will return to prep.
Notes from October, 2004 (Wright) .login script on beamppc disappeared: source /usr/local/mipp/setup-mipp-online.csh Found replacement by looking at ~alebedev/.login: source /usr/local/mipp/bin/setup_ppc.csh Switch to srt means that scripts are broken again. Found exec16 and exec24 in /usr/local/mip/releases/OLR-04.09.27/bin/Linux2.2-GCC/exec16 For now just copy them to ~/bin. I never did like srt. BaBar was cursed with this and now MIPP. initCAMAC routine is gone completely, found it in /usr/local/mipp/attic/ copied it and necessary libraries to ~/lib. Modified LD_LIBRARY_PATH (in .login) so this will work. To use new srt system (to get development executables in your path) the user must type 'srt_setup', so I put this in my .login on beamppc. beamppc:test/test_all still works, but shows that two adc modules on crate 6 are now questionable: crate=6 slot=4 0 = 111 255 crate=6 slot=6 0 = 47 63 crate=6 slot=8 0 = 109 253 should be: crate=6 slot=x 0 = 111 255 before, only slot 6 was bad. Note that pedestals look ok. Andre replied that 'crateInhibit 0' should be used instead of initCAMAC. Have to fix all scripts. Got beamppc login to work without passwords (needed to use ssh -2 protocal: cat id_rsa.pub | ssh remote 'cat - >> .ssh/authorized_keys') David report on gas 3 months of gas arrived at Fermilab (500 ft^3) cost ~$16k Terry reduced overpressure on Ckov, which resulted in reduced gas usage From Andrew Norman: Just to repeat the info, since I am assuming you didn't get it, the Jorway41 is in Crate 4, Slot 21. To dump the register use an F0 A0. The register is 12 bits wide. Be aware that while I did run a separate trigger line for you that can be controlled from this register, I didn't remove or disable your old trigger line. The CKOV should still be controlled of that original trigger line unless you modified it. I can go out and see if that is still the case if you want, but I can not imagine anyone would have changed the cabling without talking to you. The current bit assignments are: Bit 1 -- TPC 1 (dsp trig) Bit 2 -- TPC 2 (gating grid) Bit 3 -- General to Rack 5 aux Bit 4 -- CKoV In addition bits 11 and 12 are currently open, the rest have other detectors being fed off them (Let me know if you need these other assignments) The cables I ran for CKoV are labeled as "Exp Trig -> out 5" and "Exp Trig Spare 1". The drops for both are right next to the lines you had pulled. Both lines lead to the patch panel on 11b. If you have problems with either line let me know. I have no idea how the old cables are labeled, or even if they are. When I was redoing the trigger down at 11b, I just figured it was better run new (tested) good cables than to worry about tracing the old ones. You are right the information PREP has on that Jorway 41 is pretty slim. Basically they only listed the F16+ user commands. I don't have an official manual from Jorway, just some old notes that mention "F0 A0 read prog reg", your milage may vary. scan_ped script modified to take pedestals from all ADCs (takes pedestals for various widths using internal pulser). Note that some channels go down (to zero), some go up, and some more or less remain the same. 4300 manual says that for gates above 500ns need to adjust potentiometer to get non-zero pedestal. e907hv: Can not talk to hv (minicom, hv_control, hvc all fail). Device missing: /dev/ttyQ01e6 Tried reboot, did not resurrect device file. /etc/rc.d/init.d/espx restart Run this as root and you see it fails to find /usr/local/espx, which is exported from e907ana1:. Unmounted /usr/local and found espx exists in native directory. Commented out the mount in /etc/fstab and rebooted. HV now doesn't complain about /dev but still no control. Equinox box was reset, still didn't work. Ran hv_control and everything ok, probably the old feature of changing the port settings on the /dev file (see tty below). Turns out that Andre is the one who made the error in the /etc/fstab file. Reset all voltages to zero (z command in hvc), turned hv on. 10-12-04 Did pedestal scan with HV at 1000 and 1700v. No difference in pedestals compared to HV off (all data in data.oct04 subdirectory). Wrote perl script scan_ped.diff to compare the two files and look for differences. 10-13-04 Modified scan_ped to use pulser or external LED trigger. Did pedestal scan with LEDs. Find difference between LED and pulser pedestals for one channel in each ADC in crate 6. No such difference for crate 5 ADC. Also checked that pulser pedestals did not change from previous day and that pulser frequency does not matter. Checked that repeating LED pedestals results in same values, only difference is LED compared to pulser in crate 6! Modified scan_ped to scan delay. No changes observed for pedestals vs delay. (width=1000ns). 10-14-04 Wrote led_adc and ped_sub scripts to take ADC and do pedestal subtraction with LED pulser. Did HV scan and wrote led.split perl script to put output together by ADC module. Noticed strange behavior for crate 5 adc 4, which is HV slot 6. At 1500V if HV set to zero, then ADC output doesn't seem to go down. If set to 1000V, adc goes to zero. Perhaps the HV module is not really shutting off (read back says it is). Also, sometimes when set to 1500V it appears that there are no signals. This only happens for crate 5 adc 4. See problem with ADC channels. Seems like every other channel is severly attenuated. Need to check each channel with same input signal (will do at next trip). 10-17-04 Tested new adc features from David. Wrote ped_get script for generating pedestals and creating ped_set script that can download them to adcs. Wrote read_adc_ped for reading pedestal subtracted adcs (i.e. 'noinit').
10-17-04 Note that can not enable and set pedestals subtractin in 4300 when there are gets coming in! Must stop software and led pulser first. Wrote 'ped_dump' script for communicating with pedestal features of all 6 adcs. HV slot 6 @ 1500 V with LED pulser puts signals on analog line. PMT#7 has nice large signal ~500 counts above pedestal through delay board (gate 900 1200), and 1050 straight into ADC (using LLNL pigtail). Tested 3 spare 4300 boards all bad. Two read out pedestals <10 on all channels and do not register signal. Third module will not even trigger. Tested in crate 5 slot 7. Checked that the slot and cable work ok by moving ADC from crate 6 slot 4 to this slot. Worked ok, so put adc back into crate 6. Checked all channels on crate 5 slot 4 ADC with Analog #7 get 1025-1070 counts. All channels functioning ok. Crate 5 slot 4,6,8 all channels ok and pedestals all above zero for gate 900 1200. exec16 1 4 21 20 0 2048 # turn on channel 12 (2^n-1, 2^12-1 = 2048) 22 0 2048 # turn off channel 2 Implemented LED pulser remote control. Use DAC, connected inside pulser to line 7. Set to 0v for off, 6V for on. Wrote script 'led' to turn this on/off. Installed and labelled permanent cables for pulser gate for LED pulser 1 (control room) and pulser 2 (worm). Pulser gate is channel 7 of DAC in crate 6, slot 20. This gate controls both pulser outputs (Note: should only power on one pulser at a time). Exchanged modules at prep. Two Jorway 70 camac crate controllers and three lecroy 4300B ADC modules. Complained to (Mike Behnke, firstname.lastname@example.org) prep about "fermilab only" web pages for 4300 adc. Turns out that the addresses on the page were broken. Swapped out ADC crate6,slot6 with new one from prep. Bad read/write problem gone, pedestals look good. Marked old module as bad. Checked all channels of ADCs crate6, slot4,6,8. Put signal from PMT#7 directly into ADC. Everything ok. Tested other two 4300 modules from prep. Both are ok. One of the units already has terminators on it. Only checked input to a few channels (used crate6, slot 7). Some channels have pedestals at zero for gate width of 1200ns. Decided not to adjust trim pot (potentiometer "O", must remove module and module cover to change) because we will probably not have such a long gate for the real experiment anyway. Checked that pedestals >10 for 500ns width (gate 900 500): crate=5 38 43 48 61 36 33 39 49 29 44 58 42 37 37 27 42 Average 42 35 46 49 38 44 41 60 45 45 41 49 43 56 50 34 Average 76 51 83 62 77 58 75 60 75 72 81 81 62 74 21 56 Average crate=6 45 41 36 31 29 35 28 30 48 37 28 34 33 28 39 17 Average 46 32 43 35 39 43 33 33 31 29 36 34 39 36 41 26 Average 37 35 60 30 50 29 40 38 20 37 18 26 19 39 29 27 Average Checked delay board channels 49,51,53,55 and 52,54,56,58. Removed analog cable from back side of board, fed in good channel PMT#7. See proper ADC counts on all channels, so delay board and ADCs should work as advertised. Checked analog cables from detector patch panel 49-64 (bottom). From underneath detector fed good channel (PMT#63 @ 1900V) to each analog cable, proper ADC counts observed on all channels. See odd/even pattern. Odd channels all have high counts (>100) with the exception of #53. Found HV not locked in place on HV module, now all odd channels look good. Even channels barely above pedestal counts ~10. Start debugging #50. o Put analog signal onto subminex #50, get observed proper ADC counts, so base to ADC cabling ok. o Swapped HV with #51, HV tests ok for both #50 and #51. o Checked low voltage on subminex, get -5.2V ok. o Wiggled base a bit find that can change output from 2 (essentially dead) to 10 ADC counts (barely on). o Base not level with shield because someone used long and short screw to attach to detector. Replaced short screw with long screw. Replaced cable anchor that fell off during debugging. Debugging channels 65-80. Analog connections swapped at base between 71 and 72. Swapped back, now ok. Verified that HV 73 cable is bad, but it also appears that the base 73 is bad too (subminix analog and low voltage cable are good). Bottom Half View from lying underneath detector Left Beam Right | | V -----------+--------------+--------------+-----------+ 81 84 | hvslot=3 | hvslot=2 | 91 94 | | 49 53 57 61 | 65 69 73 77 | | 82 85 | 50 54 58 62 | 66 70 74 78 | 92 95 | | 51 55 59 63 | 67 71 75 79 | | 83 86 | 52 56 60 64 | 68 72 76 80 | 93 96 | +--------------+--------------+ | 87 88 hvslot=1 89 90 | -----------------------------------------------------+ 10-20-04 beamppc answers pings, but login hangs. Pushed reset button on module (abort button didn't do anything). It took at least a minute for it to reboot. Everything ok. Ran test_all. Replaced bad HV73 cable with spare. Labelled Bad and spare cables. Channel 73 appears to work, but has lower than expected ADC counts. Perhaps this is ok because channel 65 (similar one via symmetry) has unusually high ADC counts, i.e., the LED location may be responsible. Did HV scan for entire detector, 1700, 1800, 1900, 2000V. Now try TDCs. Note that this is with spider cables (which do not correspond to ADC channel order) with flip and shift, directly into 3377. Remember to run set_dac which sets thresholds to -1.8V. Note that light output from LED doesn't start until about 1.5 microseconds after beginning of pulse. gate 1400 100 seems to work the best. Don't see nice correlation with HV channel and TDC signal. Discover that turning on just a single HV channel (PMT=7 HVslot=6 1900V) causes most of TDC 1 (crate 5, slot 12) to come on: 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 0 144 180 11 510 117 111 112 106 45 113 With all HV off get occasional hits on channel 9, must be noisey. Note that TDCs are currently reading out in units of 1/2 nanosecond, so divide the above by 1/2. Put known test signal (generated from NIM discriminator in crate with LED linear fanout) into TDC via CAMAC ECL-NIM-ECL (PS Model 7126). Discriminator output is 20 nsec NIM (I wanted wider signal, but that is max module could do). Ribbon cable fro NIM-ECL to TDC. Crate 5 slot 12 all channels ok. Note that can not measure delay greater than 500 nsec (which appears as 1000 counts). Set gate to 300 nsec: 'gate 300 100'. TDC reads 725 counts. Now put digital output cable into camac ECL-NIM-ECL. Looking at NIM output see signal at 2000 nsec after LED pulser (using discriminator output, see above). See ratty signal on channel 16, see big single pulse on channel 1. Don't see any hits in TDC. Scanned delay fro 100 to 2500 nsec. Tried with and without cable flipper. Went back to direct input to TDC (with flipper). Found signals with gate set to 2300 nsec. gate 2300 100. Signals are very ratty, so get lots of multiple hits. Just take first hit. Channel map is a bit off, HV channel 1 comes out as channel 14 (in 0-15), channel 2 as 13. Switched experimental trigger line. Was 'ckov adc trig', now 'Exp Trig Out 5' (from Andrew's addressable module). 10-21-04 Fixed carriage return problem in hvc program for fast_command_send (compile on e907daq to avoid weird complaint about makedepend on e907hv). 10-22-04 Converted default scripts to use new 4300 routines (with ped subtraction capability). Tested and works ok. Created tar file backup of test directory on beamppc. Put file in my home directory on e907daq. Switched all test programs to new versions from David (2323, 3377, 4300, 5408). David modified 3377 to report in ns (before was reporting in units of 0.5 nsec) also removed reinitializing every time (takes about a second! per module!). So now '3377 1 crate slot 0' will initialize in commond stop mode, while regular read does not redo the initialization. Created init_tdc script to do the initialize all six cards. Fixed ped_set script so that 4300 are initialized properly for pedestal subtraction mode. This avoids annoying hangups. Can use gate 2300 100 for adc and tdc. The LED pulser is set long enough that there are adc counts as well, but note that these counts are not in time with TDC. Should redo relative timing study and set up double gate when at fermilab (so we can correlate TDC with ADC). Took adc and tdc data at 1900V (all.1900) and 2000V (all.2000).
11-1-04 (DMW @ FNL) beamppc is not responding again. Can ping it but not log in. Pushed restart on the front panel. This is the third time I have noticed this. First time was 10-20-04, second time was noticed by Andre last week. 11-2-04 (DMW @ FNL) CKov LED pulser 2 is laying on the floor (not in rack 17 were last seen on 10-20-04). The input gate cable is broken (conudctor was pulled out). Reinstalled pulser, replaced cable, recabelled-tied cables to rack. It looks like someone tried to remove the pulser but didn't notice the cable ties and yanked the cable out of the connector. I sent email to Holger et al asking that people let us know when stuff gets broken. test_all and read_adc (with led on) works with pulser 2 again. 11-3-04 (DMW @ FNL) Tested TDC channels 1-16 (we don't use 17-32) on all 6 modules, one at a time. All ok. Test using LED pulsers (with HV off). Put 32nsec delay (via lemo cable) in trigger (between 2323 and NIM-ECL). Use NIM output from opposite crate into NIM-ECL-NIM module in crate 5. Use ribbon cable from N-E-N to TDC to test. Move lemo cable down N-E-N and look for signal on TDC. Should see either 32ns or 16 nsec delay (depending on whether you are testing crate 5 or 6 respectively) Tested two spare 3377, labeled ok. Added feature to hvc control program (compile in e907daq). Can now set ODD or EVEN channels separately. 11-4-04 (DMW @ FNL) See if LEDs illuminate the tubes on the same side (top/bottom) or opposite side. Answer is they illuminate same side (bottom LED illuminte bottom PMTs). Notice that light output of LED goes up if you remove top/bottom, because total resistive load is reduced. 1900V, gate 1300 900 read_adc_ped LED top and bottom flashing 212 11 539 14 22 1 2047 48 2 13 1299 17 326 14 1288 2 Average 49 0 317 32 1280 2 2047 1 276 14 1508 18 619 13 427 0 Average 2047 7 239 119 8 582 2047 190 100 45 91 11 12 2 11 33 Average 135 3 588 12 580 17 379 26 149 12 461 15 413 15 1137 6 Average 2047 33 996 20 1531 30 1607 20 35 3 781 2 624 21 1063 5 Average 133 3 78 114 11 122 142 36 156 215 482 6 27 825 5 13 Average LED top flashing only (unplug bottom) 435 12 1089 21 39 1 2047 81 3 10 2047 28 670 15 2047 4 Average 80 0 569 53 2047 1 2047 6 654 22 2047 35 1354 16 924 0 Average 2047 7 357 249 14 1089 2047 394 174 103 156 18 17 2 18 47 Average 6 1 2 2 15 10 0 4 16 8 1 2 14 9 6 2 Average 25 14 3 0 12 15 3 0 1 2 0 1 10 1 1 2 Average 4 2 1 4 1 2 1 7 1 5 8 0 2 2 1 0 Average LED top flashing only (unplug top) 20 7 1 1 1 0 4 2 0 9 12 1 17 9 4 0 Average 16 0 2 1 11 8 4 0 9 7 4 0 12 9 1 0 Average 1 0 1 11 1 0 2 2 1 0 8 1 1 0 1 0 Average 291 2 1409 23 1146 18 764 50 348 18 1181 26 825 18 2047 12 Average 2047 28 1907 9 2047 40 2047 33 63 6 1902 6 1340 23 2045 11 Average 196 6 120 253 15 211 277 91 303 482 1118 11 60 1236 5 17 Average LED top and bottom flashing 200 11 561 13 22 1 2047 49 2 11 1310 17 329 14 1279 2 Average 46 1 293 29 1275 0 2047 0 271 17 1520 18 617 13 416 0 Average 2047 7 240 114 8 364 2047 193 101 45 108 10 10 1 11 33 Average 130 4 589 11 575 17 380 26 145 12 457 14 414 12 1148 6 Average 2047 6 1010 20 1534 29 1580 19 36 3 797 2 623 19 1084 5 Average 134 3 77 116 10 123 137 35 155 211 492 5 30 831 5 12 Average OK, Now try to increase the amount of light seen by the odd (relative) numbered PMTs. Turn on a single card and turn on just the odd channel HV. Tried to raise the output voltage, but the linear fanout maxes out at pulser voltage setting of about 5.5, was already running at 5V (pulser 2 which is in the worm). So can't get any more amplitude. hvslot=3 crate=6 ADC=4 PMTS=49-64 ODD=1900V Even=0V > 4300 1 6 4 100 noinit | grep Average 0 2 1 9 1 15 1 22 1 9 1 10 1 10 1 5 Average Increase the LED pulse width by factor of 10 to 20 microseconds width (normally is 2) and increase the ADC gate: ped_get 1300 3000. Note that ped_get will set the width delay and then execute ped_set. Raise HV to 2000V for ODD, still 0 for EVEN. > led on ; 4300 1 6 4 100 noinit | grep Average turning LED pulser on 0 1 0 152 0 177 1 315 1 148 1 119 0 195 1 41 Average Lots of light now, so turn on ODD channels of hvslot=2 crate=6 ADC=6: 0 0 0 152 1 176 1 330 2 150 2 121 0 195 1 39 Average 7 137 5 222 5 319 5 148 7 4 5 4 5 126 2 36 Average Wrote hv script on e907hv:hvc to control hv from command line Wrote plateau script to plateau a single hv slot. Wrote plateau.all script to plateau all adc slots. Wrote plateau.each script to plateau each channel. From adc*.plateau files: adc3 2047 45 871 2047 116 12 2047 2047 689 1177 1306 85 46 8 0 116 adc2 1500 25 2047 349 2047 243 2047 139 2047 233 2047 87 2047 170 2047 42 adc1 2047 183 2047 136 148 22 2047 437 2 234 2047 187 2047 187 2047 2 adc4 2047 32 2047 180 2047 193 2047 323 2047 198 2047 142 2047 228 2047 81 adc5 2047 240 2047 270 2047 332 2047 132 536 6 2047 8 2047 137 2047 41 adc6 647 19 387 2047 88 1623 2047 1711 2047 2047 2047 64 338 2047 19 74 adc3 dead? dead?dead adc2 adc1 dead dead adc4 adc5 dead dead adc6 dead It turns out there is significant cross talk between channels, especially when the a channel saturates the adc. Refer to the individual channel plateaus (on e907hv) to see which channels have no signal. no signal channels (0:15) adc1: 8, 15 adc2: 1, 15 adc3: 5, 13, 14 adc4: 1 adc5: 9, 11 adc6: 1, 15 channel 14 pmt is connected to #15 readout! ADC to PMT map 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 Top (Crate 5) 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 adc1 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 adc2 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 adc3 Bottom (Crate 6) 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 adc4 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 adc5 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 adc6 1st Check all top PMT that do not see light Check adc6 pmt 14/15 swap? Check weak channel output reduced by lowering LED voltage adc1: 15 fixed, delay line chip was loose adc1: 8 signal reaches delay board, fixed loose wire from delay chip but also lemo connector is broken on user side! must resolder adc2: 1 & 15 patch panel to adc is ok, no signal at patch panel 11-9-04 beamppc has problem with ssh again, cannot get plateau scripts on e907hv to work anymore, beamppc hangs on ssh. Stopped DAQ (was running) and rebooted beamppc. Worked for about 3 ssh sessions, then hung up again! adc2: 1 analog line from base is ok hv is ok Low voltage power is bad, it reads -5.2V, but base doesn't work base does work if use another low voltage cable adc2: 11 ? redo plateau, check at constant HV because ADC counts seem to drop with time adc3:14 hv ok, analog from patch panel ok adc3: 10 for a while the output was much lower than last week, now ok adc=1187 @ 1900, earlier this evening it was only 12! adc6: 14 HV cables are labelled correctly analog lines from patch panel to delay board are labelled corretly subminex to patch panel cables are mislablled change labels and connect correctly. now 14 is ok. 11-10-04 adc6: 15 HV, analog, lv ok adc6: 1 HV, analog, lv ok 4 was swapping lv with 1 and now after putting lv back the output is fluctuating, was consistently about 40 but now is often zero! adc6: 7 pedestal has changed dramatically now 130 counts higher. redid ped_get 1300 3000 and now seems ok, see light on it. adc4: 1(50) analog, HV, lv ok adc5: b9(74) analog ok, HV BAD adc5: 11(76) base very wobbly, analog, hv, lv ok ssh to beamppc is horribly slow, it effectively kills plateau scripts. ssh -v says it is the exec after authentication that is the problem -x doesn't help, but switching to -1 does! modified plateau scripts and we are back in business. adc6: 4(85) redo plateau, is about factor of two lower now, but still see signal adc2: 11(28) redid plateau and it looks fine (e907hv:adc_data3) but ADC drops with time to zero! swapped hv with another channel, hv is fine, it is the base/tube adc3: 14(47) replaced analog line between base and patch panel, now ok Turned off low voltage to check subminex cable, power back on and pedestal for adc3:14(45) is now lower by 25 counts. Or maybe just because cable was bad (see above) and I didn't notice after changing it. Redo pedestals: ped_get 1300 3000. adc2: 1(18) need to replace low voltage subminex Made low voltage splitter box. Used low voltage from mirror 41, split to 41 and 18. 11-11-04 Crash during afs editing of this file. Lost all notes from today! Reentering some of it. Will work locally now. adc1: 8(9) replaced lemo on delay board, ok adc5: 9(74) replaced bad HV cable, ok Check TDCs. With HV and LV off get few noisey channels. With LV on get lots fo noisey channels LV off. 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 12 0 414 14 16 12 0 141 22 14 0 252 16 set_dac -1.8, LV on 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 12 0 448 465 14 0 307 16 0 0 0 0 110 46 189 98 279 12 0 333 147 221 14 0 81 251 9 32 476 312 428 89 16 set_dac -1 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 12 0 331 5 219 88 118 160 14 0 33 193 20 76 15 16 0 39 189 12 0 131 217 68 259 36 416 330 219 14 0 68 373 64 463 389 105 208 361 198 16 0 44 18 19 18 498 26 62 69 set dac -3 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 12 0 229 419 14 16 0 181 293 166 12 0 156 14 0 120 288 105 120 419 321 394 129 16 How to restart ifix server (from Holger) PC in Mtest is pixel06. Click Login/Logout Logout Public Login Holger, killi Ctrl-W to get ifix window such that you can minimize it, so can get to Windows Start Menu. Start:Run, cmd cd hmeyer iFix2XmlRpc 80 Minimize Ctrl-W Logout Holger Login Public,
Nominal height of Cerenkov box is 333cm from top of box to bottom of bracket into which adjustable foot is attached. The foot adds an additional ~4cm
This was measured with tape measure by DMW DJL, Jan 2002.
Extra space is required on both sides of box versus nominal dimensions given in Ed's memory dump.